------------------------------------------------------------ Date: Thu Feb 19 08:10:20 2004, Shift: 00:00-08:00 Supervisor: RK, with: DO Log book#15 pages 61 - 61 Shift Summary: ------------------------------------------------------------ 1. In spite of AtR problems they were able to inject at 1.30 and keep the store for more than three hours, when blue quenched. Since 6am there is Controlled Access. 2. Pedestals 10417 and 10420 (IO- and Run-servers problems caused 2 runs missing) Data runs: 10421 - FS@4deg@A3450 tr.2 132k, MRS@45deg@B1050 tr.3 150k 10422 - FS@4deg@A3450 tr.2 82k, MRS@45deg@B 500 tr.3 120k 10423 - FS@4deg@A3450 tr.2 67k, MRS@45deg@A 500 tr.3 120k 10424 - FS@4deg@A3450 tr.2 45k, MRS@45deg@A 700 tr.3 81k 3. No detector problems, however: - new magnet software does not have possibility to set other fields than those predefined in program, thus in order to make runs with D5 at 500 I had to use old software. I guess it is possible to add some other settings. - number of mice in the CH is growing... also thanks to the chocolate bars (I think they like Nestle's Treasures with peanut butter). So please do check for small bites before eating chocolates. ------------------------------------------------------------ _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Thu Feb 19 08:11:18 2004
This archive was generated by hypermail 2.1.8 : Thu Feb 19 2004 - 08:11:39 EST